# On the Robustness of Kernel Goodness-of-Fit Tests

### v3

@@ 5.1.4 Misspecified Tails
Finally, we now change the data-generating mechanism to study the impact of misspecified tails. We sample from $Q_{\nu}=t_{\nu}\sqrt{(\nu-2)/\nu}$, where $t_{\nu}$ is the Student’s t-distribution with degree-of-freedom (dof) $\nu>2$, and the scaling factor $\sqrt{(\nu-2)/\nu}$ ensures its second moment matches that of $P$. We set the uncertainty radius so that the R-KSD test is robust to $Q_{\nu}$ with $\nu\geq\nu_{0}$ for different values of $\nu_{0}$. This is achieved using Proposition 6 to derive a bound on $D(Q_{\nu},P)$, which is discussed in detail in Proposition 22 of Section A.6.4. The results are reported in Figure 4. For small values of $\nu$, the tail of the model is misspecified, leading the robust tests with $\nu_{0}=10$ or $20$ to reject with high probability because the KSD between $Q_{\nu}$ and $P$ is large. As $\nu$ grows, $Q_{\nu}$ converges weakly to the standard Gaussian, making the model well-specified in the limit $\nu\to\infty$. Consequently, neither the standard tests or the robust tests reject the null hypothesis for large $\nu$. The robust test with $\nu_{0}=5$ shows no power because, to be robust to tails that are so misspecified, the uncertainty radius $\theta$ computed using Proposition 6 must be very large, thus resulting in a small test statistic (cf. (9)) and low test power. Overall, the robust tests are more conservative than the standard one. This is because the robust tests control all possible contaminations within a KSD-ball, while misspecified tails are only one of the many possible such forms of contamination.
@@ 5.2 Gaussian-Bernoulli Restricted Boltzmann Machines
Our next experiment is an example of energy-based model called a Gaussian-Bernoulli Restricted Boltzmann Machine (RBM) model (Cho et al.,, 2013). RBMs have been used in a wide range of scientific applications (see Fischer and Igel,, 2014; Zhang et al.,, 2019, for a review), and they are a common benchmarks for assessing kernel GOF tests (Liu et al.,, 2016; Jitkrittum et al.,, 2017; Schrab et al.,, 2022). RBMs are latent-variable models with joint density $p(\mathbf{x},\mathbf{h})\propto\exp(\frac{1}{2}\mathbf{x}^{\top}B\mathbf{h}+\mathbf{b}^{\top}\mathbf{x}+\mathbf{c}^{\top}\mathbf{h}-\frac{1}{2}\|\mathbf{x}\|_{2}^{2})$, where $\mathbf{x}\in\mathbb{R}^{d}$ is an observable variable, $\mathbf{h}\in\{\pm 1\}^{d^{\prime}}$ is a binary hidden variable with latent dimension $d^{\prime}$, and $B\in\mathbb{R}^{d\times d^{\prime}},\mathbf{b}\in\mathbb{R}^{d}$ and $\mathbf{c}\in\mathbb{R}^{d^{\prime}}$ are model parameters. Computing the normalizing constant of the marginal $p(\mathbf{x})$ requires summing over $2^{d^{\prime}}$ terms, so it becomes intractable when $d^{\prime}$ is large. However, its score function has a closed-form expression $\mathbf{s}_{p}(\mathbf{x})=b-\mathbf{x}+B\mathrm{tanh}(B^{\top}\mathbf{x}+\mathbf{c})$, where $\mathrm{tanh}$ is applied entry-wise. We set $d=50$ and $d^{\prime}=10$, and randomly initialize $B,\mathbf{b},\mathbf{c}$ by sampling each entry independently from a $\mathcal{N}(0,1)$. We then generate $n=500$ samples from the model by block Gibbs sampling following Cho et al., (2013); Jitkrittum et al., (2017), and randomly replace $\epsilon\in[0,1]$ proportions of the data with outliers drawn from $R=\mathcal{N}(0,0.1^{2}I_{d})$; see the left and middle of Figure 5. As shown by the plots, the outliers are clearly abnormal since they have low density values.
@@
We plot the rejection probabilities of the tests on the right-hand side of Figure 5. The standard IMQ-KSD and Tilted-KSD tests are not well-calibrated under the composite null $H^{\mathrm{C}}_{0}:Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$, while robust-KSD with $\theta$ chosen by setting $\epsilon_{0}=0.1$ controls the rejection probability below the level $\alpha=0.05$ when the contamination ratio $\epsilon$ does not exceed $\epsilon_{0}$. On the other hand, when $\epsilon$ exceeds the maximal allowed proportion $\epsilon_{0}$, the robust test rejects $H^{\mathrm{C}}_{0}$ with probability approaching $1$, thus showing its power.
@@ 5.3 A Kernel Exponential Family Model for Density Estimation
The results are shown in Figure 6. We first study the level of our tests in Figure 6(a). As the outlier mean $z$ increases, the IMQ-KSD test rejects the null $H_{0}:Q=P$ with increasing probability, suggesting that this is due to the presence of outliers. Notably, this happens even though the parameter estimator is not susceptible to contamination; see the bottom row of Figure 6(a) for the fitted model with and without outliers. When the tilted kernel was used, the standard KSD test has a lower rejection probability, although it is still above the nominal level. On the other hand, the robust-KSD test is well-calibrated for all values of $z$.
We then study the power of KSD tests in Figure 6(b). To show the robust test has non-trivial power when no outliers are present, we fit the model parameters $\hat{\eta}$ on half of the data (without contamination) and replace the first component by $\hat{\eta}_{1}+\delta$, where $\delta\in\{0,0.5,1,2\}$ is an error offset, so that the resulting model has a poor fit. Other experimental setups including the choice of $\theta$ remain the same as before. As expected, the robust test has lower power than the standard tests; however, as the offset value $\delta$ increases, the robust test eventually rejects with high probability, thus showing its power under no contamination.
@@ 5.4 Limitations for Multimodal Models
The results are reported in Figure 7. As the scale $\gamma$ increases, the modes become more separated, leading to smaller values of $D(Q,P)$, as shown in the left and middle plots. Consequently, $Q$ eventually becomes indistinguishable from $P$ under KSD, and all tests fail to reject with high probability. This is a result of the aforementioned myopia of KSD, which affects both the standard and the robust tests and regardless of whether IMQ or tilted kernels are used. The opposite is observed when $\gamma$ is small and the modes overlap, where all tests can correctly reject the null with high probability. Compared with the standard tests, the robust one is more conservative because it controls all possible contamination; nevertheless, it still achieves non-trivial power when $\epsilon_{0}$, thus also the radius $\theta$, is set to a small value.
@@ 6 Conclusion and Discussion
Secondly, as discussed in Section 2.3, our proposed robust KSD test does not require the Stein kernel $u_{p}$ to be bounded. A bounded Stein kernel is convenient as the resulting KSD-ball $\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ contains many contamination models of interest, such as Huber’s model with an arbitrary noise distribution. However, this could be too strict in some applications where outliers cannot take an arbitrary value, e.g., when the domain is bounded or the data acquisition procedure is highly regulated to prevent extreme outliers. In these cases, even with an unbounded kernel, the KSD-ball could still capture all contamination models of interest. Unbounded Stein kernels possess appealing theoretical properties such as convergence-control of moments (Kanagawa et al.,, 2022). Exploring whether such kernels can lead to more powerful robust-KSD tests is another worthwhile future research direction.
@@ Appendix A Proofs of Theoretical Results
A.1 Proof of Theorem 1 and Related Preliminary Results
Overview of proof. For any $\epsilon\in[0,1]$ and $R\in\mathcal{P}(\mathbb{R}^{d})$, a random variable $\mathbf{X}_{i}$ drawn from $Q=(1-\epsilon)P+\epsilon R$ can be written as $\mathbf{X}_{i}=(1-\xi_{i})\mathbf{X}_{i}^{\ast}+\xi_{i}\mathbf{Z}_{i}$, where $\xi_{i}\sim\mathrm{Bernoulli}(\epsilon)$, $\mathbf{X}_{i}^{\ast}\sim P$ and $\mathbf{Z}_{i}\sim R$ are independent, and the equality is understood as equality in distribution. Defining the random variable $M^{\prime}=\sum_{i=1}^{n}\xi_{i}$, then $M^{\prime}\sim\mathrm{Binomial}(n,\epsilon)$ and represents the number of contaminated data in a random sample $\mathbb{X}_{n}$ drawn from $Q$. We will first show that the conclusion in Theorem 1 holds conditional on the event $\{M^{\prime}\geq m^{\prime}_{0}\}$ for some integer $m^{\prime}_{0}\in[0,n]$, i.e., when there are at least $m^{\prime}$ contaminated data in the sample. This will be sufficient to show Theorem 1 by proving that this event occurs with high probability.
@@ A.1 Proof of Theorem 1 and Related Preliminary Results
  • We first present three preliminary results. Lemma 7 and Lemma 8 respectively bound the test statistic $D^{2}(\mathbb{X}_{n,\mathbf{z}})$ and $q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n,\mathbf{z}})$ for a given number of corrupted data, $m^{\prime}$, by quantities of the form $a_{n}u_{p}(\mathbf{z},\mathbf{z})+b_{n}$ for some $a_{n},b_{n}$. These results allow us to compare the relative rate at which $D^{2}(\mathbb{X}_{n,\mathbf{z}})$ and $q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n,\mathbf{z}})$ scale with $\mathbf{z}$. Both results will be used to prove Proposition 9, which states that the claim in Theorem 1 holds conditionally on the number of corrupted data.
  • Section A.1.1, A.1.2, A.1.3 and A.1.4 provide the proofs for Lemma 7, Lemma 8, Proposition 9 and Theorem 1, respectively.
@@
Assume $\sigma_{\infty}^{2}\coloneqq\sup_{\mathbf{z}\in\mathbb{R}^{d}}\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\mathbf{z})^{2}]<\infty$ and suppose 1 holds. Then, for any $\delta>0$ and integers $0<m^{\prime}<n$, the following event holds with probability at least $1-\delta$:
@@
Suppose 1 holds. Assume (i) $\sigma_{\infty}^{2}\coloneqq\sup_{\mathbf{z}\in\mathbb{R}^{d}}\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\mathbf{z})^{2}]<\infty$ and (ii) $\xi_{P}^{2}\coloneqq\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\mathbf{X})^{2}]<\infty$. Then there exist constants $C,C_{P,1}$, and $C_{P,2}$ depending only on $P$ such that, for any $\delta>0$ and integers $0<m^{\prime}<n$, the following event holds with probability at least $1-\delta$:
@@
There exists a constant $C>0$ such that the following holds. Suppose 1 holds. Also assume the following integrability conditions
@@
[…]
@@
[…]
@@ A.1.1 Proof of Lemma 7
Bounding the second term in (10) requires bounding $S_{m,\mathbf{z}}$. Because by assumption $\mathbf{X}_{i}=\mathbf{X}_{i}^{\ast}\sim P$ are i.i.d. for $i=1,\ldots,m$, the term $S_{m,\mathbf{z}}$ is a sum of i.i.d. random variables $u_{p}(\mathbf{X}_{i}^{\ast},\mathbf{z})$. Moreover, $u_{p}(\mathbf{X}_{i}^{\ast},\mathbf{z})$ are zero-mean, because the proof of Gorham and Mackey, (2017, Proposition 1) shows that $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[u_{p}(\mathbf{X}^{\ast},\cdot)]=0$ whenever $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[\|\mathbf{s}_{p}(\mathbf{X}^{\ast})\|_{2}]<\infty$, which holds under 1. We can therefore compute the variance of $S_{m,\mathbf{z}}$ as

### v4

@@ 5.1.4 Misspecified Tails
Finally, we now change the data-generating mechanism to study the impact of misspecified tails. We sample from $Q_{\nu}=t_{\nu}\sqrt{(\nu-2)/\nu}$, where $t_{\nu}$ is the Student’s t-distribution with degree-of-freedom (dof) $\nu>2$, and the scaling factor $\sqrt{(\nu-2)/\nu}$ ensures its second moment matches that of $P$. We set the uncertainty radius so that the R-KSD test is robust to $Q_{\nu}$ with $\nu\geq\nu_{0}$ for different values of $\nu_{0}$. This is achieved using Proposition˜6 to derive a bound on $D(Q_{\nu},P)$, which is discussed in detail in Proposition˜22 of Section˜A.6.4. The results are reported in Figure˜4. For small values of $\nu$, the tail of the model is misspecified, leading the robust tests with $\nu_{0}=10$ or $20$ to reject with high probability because the KSD between $Q_{\nu}$ and $P$ is large. As $\nu$ grows, $Q_{\nu}$ converges weakly to the standard Gaussian, making the model well-specified in the limit $\nu\to\infty$. Consequently, neither the standard tests or the robust tests reject the null hypothesis for large $\nu$. The robust test with $\nu_{0}=5$ shows no power because, to be robust to tails that are so misspecified, the uncertainty radius $\theta$ computed using Proposition˜6 must be very large, thus resulting in a small test statistic (cf. (9)) and low test power. Overall, the robust tests are more conservative than the standard one. This is because the robust tests control all possible contaminations within a KSD-ball, while misspecified tails are only one of the many possible such forms of contamination.
@@ 5.2 Gaussian-Bernoulli Restricted Boltzmann Machines
Our next experiment is an example of energy-based model called a Gaussian-Bernoulli Restricted Boltzmann Machine (RBM) model (Cho et al.,, 2013). RBMs have been used in a wide range of scientific applications (see Fischer and Igel,, 2014; Zhang et al.,, 2019, for a review), and they are a common benchmarks for assessing kernel GOF tests (Liu et al.,, 2016; Jitkrittum et al.,, 2017; Schrab et al.,, 2022). RBMs are latent-variable models with joint density $p(\mathbf{x},\mathbf{h})\propto\exp(\frac{1}{2}\mathbf{x}^{\top}B\mathbf{h}+\mathbf{b}^{\top}\mathbf{x}+\mathbf{c}^{\top}\mathbf{h}-\frac{1}{2}\|\mathbf{x}\|_{2}^{2})$, where $\mathbf{x}\in\mathbb{R}^{d}$ is an observable variable, $\mathbf{h}\in\{\pm 1\}^{d^{\prime}}$ is a binary hidden variable with latent dimension $d^{\prime}$, and $B\in\mathbb{R}^{d\times d^{\prime}},\mathbf{b}\in\mathbb{R}^{d}$ and $\mathbf{c}\in\mathbb{R}^{d^{\prime}}$ are model parameters. Computing the normalizing constant of the marginal $p(\mathbf{x})$ requires summing over $2^{d^{\prime}}$ terms, so it becomes intractable when $d^{\prime}$ is large. However, its score function has a closed-form expression $\mathbf{s}_{p}(\mathbf{x})=b-\mathbf{x}+B\mathrm{tanh}(B^{\top}\mathbf{x}+\mathbf{c})$, where $\mathrm{tanh}$ is applied entry-wise. We set $d=50$ and $d^{\prime}=10$, and randomly initialize $B,\mathbf{b},\mathbf{c}$ by sampling each entry independently from a $\mathcal{N}(0,1)$. We then generate $n=500$ samples from the model by block Gibbs sampling following Cho et al., (2013); Jitkrittum et al., (2017), and randomly replace $\epsilon\in[0,1]$ proportions of the data with outliers drawn from $R=\mathcal{N}(0,0.1^{2}I_{d})$; see the left and middle of Figure˜5. As shown by the plots, the outliers are clearly abnormal since they have low density values.
@@
We plot the rejection probabilities of the tests on the right-hand side of Figure˜5. The standard IMQ-KSD and Tilted-KSD tests are not well-calibrated under the composite null $H^{\mathrm{C}}_{0}:Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$, while robust-KSD with $\theta$ chosen by setting $\epsilon_{0}=0.1$ controls the rejection probability below the level $\alpha=0.05$ when the contamination ratio $\epsilon$ does not exceed $\epsilon_{0}$. On the other hand, when $\epsilon$ exceeds the maximal allowed proportion $\epsilon_{0}$, the robust test rejects $H^{\mathrm{C}}_{0}$ with probability approaching $1$, thus showing its power.
@@ 5.3 A Kernel Exponential Family Model for Density Estimation
The results are shown in Figure˜6. We first study the level of our tests in Figure˜6(a). As the outlier mean $z$ increases, the IMQ-KSD test rejects the null $H_{0}:Q=P$ with increasing probability, suggesting that this is due to the presence of outliers. Notably, this happens even though the parameter estimator is not susceptible to contamination; see the bottom row of Figure˜6(a) for the fitted model with and without outliers. When the tilted kernel was used, the standard KSD test has a lower rejection probability, although it is still above the nominal level. On the other hand, the robust-KSD test is well-calibrated for all values of $z$.
We then study the power of KSD tests in Figure˜6(b). To show the robust test has non-trivial power when no outliers are present, we fit the model parameters $\hat{\eta}$ on half of the data (without contamination) and replace the first component by $\hat{\eta}_{1}+\delta$, where $\delta\in\{0,0.5,1,2\}$ is an error offset, so that the resulting model has a poor fit. Other experimental setups including the choice of $\theta$ remain the same as before. As expected, the robust test has lower power than the standard tests; however, as the offset value $\delta$ increases, the robust test eventually rejects with high probability, thus showing its power under no contamination.
@@ 5.4 Limitations for Multimodal Models
The results are reported in Figure˜7. As the scale $\gamma$ increases, the modes become more separated, leading to smaller values of $D(Q,P)$, as shown in the left and middle plots. Consequently, $Q$ eventually becomes indistinguishable from $P$ under KSD, and all tests fail to reject with high probability. This is a result of the aforementioned myopia of KSD, which affects both the standard and the robust tests and regardless of whether IMQ or tilted kernels are used. The opposite is observed when $\gamma$ is small and the modes overlap, where all tests can correctly reject the null with high probability. Compared with the standard tests, the robust one is more conservative because it controls all possible contamination; nevertheless, it still achieves non-trivial power when $\epsilon_{0}$, thus also the radius $\theta$, is set to a small value.
@@ 6 Conclusion and Discussion
Secondly, as discussed in Section˜2.3, our proposed robust KSD test does not require the Stein kernel $u_{p}$ to be bounded. A bounded Stein kernel is convenient as the resulting KSD-ball $\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ contains many contamination models of interest, such as Huber’s model with an arbitrary noise distribution. However, this could be too strict in some applications where outliers cannot take an arbitrary value, e.g., when the domain is bounded or the data acquisition procedure is highly regulated to prevent extreme outliers. In these cases, even with an unbounded kernel, the KSD-ball could still capture all contamination models of interest. Unbounded Stein kernels possess appealing theoretical properties such as convergence-control of moments (Kanagawa et al.,, 2022). Exploring whether such kernels can lead to more powerful robust-KSD tests is another worthwhile future research direction.
@@ Appendix A Proofs of Theoretical Results
A.1 Proof of Theorem˜1 and Related Preliminary Results
Overview of proof. For any $\epsilon\in[0,1]$ and $R\in\mathcal{P}(\mathbb{R}^{d})$, a random variable $\mathbf{X}_{i}$ drawn from $Q=(1-\epsilon)P+\epsilon R$ can be written as $\mathbf{X}_{i}=(1-\xi_{i})\mathbf{X}_{i}^{\ast}+\xi_{i}\mathbf{Z}_{i}$, where $\xi_{i}\sim\mathrm{Bernoulli}(\epsilon)$, $\mathbf{X}_{i}^{\ast}\sim P$ and $\mathbf{Z}_{i}\sim R$ are independent, and the equality is understood as equality in distribution. Defining the random variable $M^{\prime}=\sum_{i=1}^{n}\xi_{i}$, then $M^{\prime}\sim\mathrm{Binomial}(n,\epsilon)$ and represents the number of contaminated data in a random sample $\mathbb{X}_{n}$ drawn from $Q$. We will first show that the conclusion in Theorem˜1 holds conditional on the event $\{M^{\prime}\geq m^{\prime}_{0}\}$ for some integer $m^{\prime}_{0}\in[0,n]$, i.e., when there are at least $m^{\prime}$ contaminated data in the sample. This will be sufficient to show Theorem˜1 by proving that this event occurs with high probability.
@@ A.1 Proof of Theorem˜1 and Related Preliminary Results
  • We first present three preliminary results. Lemma˜7 and Lemma˜8 respectively bound the test statistic $D^{2}(\mathbb{X}_{n,\mathbf{z}})$ and $q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n,\mathbf{z}})$ for a given number of corrupted data, $m^{\prime}$, by quantities of the form $a_{n}u_{p}(\mathbf{z},\mathbf{z})+b_{n}$ for some $a_{n},b_{n}$. These results allow us to compare the relative rate at which $D^{2}(\mathbb{X}_{n,\mathbf{z}})$ and $q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n,\mathbf{z}})$ scale with $\mathbf{z}$. Both results will be used to prove Proposition˜9, which states that the claim in Theorem˜1 holds conditionally on the number of corrupted data.
  • Section˜A.1.1, A.1.2, A.1.3 and A.1.4 provide the proofs for Lemma˜7, Lemma˜8, Proposition˜9 and Theorem˜1, respectively.
@@
Assume $\sigma_{\infty}^{2}\coloneqq\sup_{\mathbf{z}\in\mathbb{R}^{d}}\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\mathbf{z})^{2}]<\infty$ and suppose ˜1 holds. Then, for any $\delta>0$ and integers $0<m^{\prime}<n$, the following event holds with probability at least $1-\delta$:
@@
Suppose ˜1 holds. Assume (i) $\sigma_{\infty}^{2}\coloneqq\sup_{\mathbf{z}\in\mathbb{R}^{d}}\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\mathbf{z})^{2}]<\infty$ and (ii) $\xi_{P}^{2}\coloneqq\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\mathbf{X})^{2}]<\infty$. Then there exist constants $C,C_{P,1}$, and $C_{P,2}$ depending only on $P$ such that, for any $\delta>0$ and integers $0<m^{\prime}<n$, the following event holds with probability at least $1-\delta$:
@@
There exists a constant $C>0$ such that the following holds. Suppose ˜1 holds. Also assume the following integrability conditions
@@
$n^{\frac{(1-s)}{2}}\;>\;2\sqrt{2}C\log\left(\frac{4}{\alpha}\right)$
@@
$\Pr\nolimits_{\mathbb{X}_{n-m^{\prime}}^{\ast}\sim P,\mathbf{W}}\big{(}D^{2}(\mathbb{X}_{n,\mathbf{z}})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n,\mathbf{z}})\big{)}\;\geq\;1-n^{-(1-s)}.$
@@ A.1.1 Proof of Lemma˜7
Bounding the second term in (10) requires bounding $S_{m,\mathbf{z}}$. Because by assumption $\mathbf{X}_{i}=\mathbf{X}_{i}^{\ast}\sim P$ are i.i.d. for $i=1,\ldots,m$, the term $S_{m,\mathbf{z}}$ is a sum of i.i.d. random variables $u_{p}(\mathbf{X}_{i}^{\ast},\mathbf{z})$. Moreover, $u_{p}(\mathbf{X}_{i}^{\ast},\mathbf{z})$ are zero-mean, because the proof of Gorham and Mackey, (2017, Proposition 1) shows that $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[u_{p}(\mathbf{X}^{\ast},\cdot)]=0$ whenever $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[\|\mathbf{s}_{p}(\mathbf{X}^{\ast})\|_{2}]<\infty$, which holds under ˜1. We can therefore compute the variance of $S_{m,\mathbf{z}}$ as